# FPC-VLA: A Vision-Language-Action Framework with a Supervisor for Failure Prediction and Correction

### v1

@@ IV Experiment
RT-1 [5] 85.7 44.2 73.0 6.5 52.4 89.8 50.0 32.3 2.6 43.7
RT-1-X [7] 56.7 31.7 59.7 21.3 42.4 49.0 32.3 29.4 10.1 30.2
RT-2-X [7] 78.7 77.9 25.0 3.7 46.3 82.3 79.2 35.3 20.6 54.4
Octo-Base [8] 17.0 4.2 22.7 0.0 11.0 0.6 3.1 1.1 0.0 1.2
OpenVLA [9] 18.0 56.3 63.0 0.0 34.3 60.8 67.7 28.8 0.0 39.3
RoboVLMs [12] 77.3 61.7 43.5 24.1 41.9
CogACT [13] 91.3 80.5 71.8 50.9 74.8 89.6 80.8 28.3 46.6 61.3
SpatialVLA [10]
@@
TABLE V: Comparison results on LIBERO with Franka
@@
Diffusion [28] 68.3 78.3 82.5 50.5 72.4
Octo-Base [8] 84.6 78.9 85.7 51.1 75.1
OpenVLA [9] 79.2 84.7 88.4 53.7 76.5
TravceVLA [11] 75.1 84.6 85.2 54.1 74.8
SpatialVLA [10] 78.6 88.2 89.9 55.5 78.1
ThinkAct [27] 87.1 88.3 91.4 70.9 84.4
Cot-VLA [29] 87.6 87.5 91.6 69.0 81.1
@@ IV-A Implementation Details
The VLM model of FPC-VLA adopts Prismatic VLMs [30]. The VLM model of the supervisor uses Qwen2.5-vl 7B [31]. During training, we use 8 NVIDIA H100 GPUs. For the VLA model, the batch size is set to 256, the learning rate is $2\times 10^{-5}$, and 15 action steps are predicted per inference. For the supervisor model, training is conducted with bfloat16 precision, a batch size of 64, a learning rate of $10^{-4}$, LoRA rank of 8, LoRA alpha of 32, and max pixels of 1003520. During inference, for WidowX, the supervisor is trained on dataset generated from BridgeV2 [25] (408771 entries), with $\delta_{\text{trans}}^{\text{small}},\delta_{\text{trans}}^{\text{large}},\delta_{\text{rot}}^{\text{small}},\delta_{\text{rot}}^{\text{large}}$ set to 0.001, 0.01, 0.001, and 0.01 respectively, and the size of action window $N=7$. For Google Robot, dataset is generated from [5] (728760 entries), with $\delta_{\text{trans}}^{\text{small}},\delta_{\text{trans}}^{\text{large}},\delta_{\text{rot}}^{\text{small}},\delta_{\text{rot}}^{\text{large}}$ set to 0.01, 0.1, 0.01, and 0.1 respectively, and $N=2$. For Franka, dataset is generated from [26] (15604 entries), with $\delta_{\text{trans}}^{\text{small}},\delta_{\text{trans}}^{\text{large}},\delta_{\text{rot}}^{\text{small}},\delta_{\text{rot}}^{\text{large}}$ set to 0.01, 0.1, 0.001, and 0.01 respectively. $N=5$ for the goal setting, $N=7$ for the spatial setting, and $N=11$ for others. In action fusion module, $\alpha=0.1$, $\lambda=\beta=0.01$. Supervisor is called no more than 3 times in each task.
FPC-VLA achieves inference time of 0.176s for non-keyframes and 1.766s for keyframes. As the supervisor is called at most three times during the hundreds of inferences per task, its benefit becomes more evident in long-horizon tasks. For example, in the Open Top Drawer and Place Apple task, the total inference time of FPC-VLA is 40.5 s, which is only slightly higher than the 35.2 s when the supervisor is inactive.
IV-B Evaluation on Simulation Benchmarks
@@
To evaluate the FPC-VLA model, we conduct experiments on two simulation platforms—SIMPLER [32] and LIBERO [26] —as well as on three types of robotic arms: WidowX, Google Robot, and Franka. On SIMPLER, we primarily highlight FPC-VLA’s zero-shot learning capability, using only pre-training on the OXE dataset [7] without domain-specific fine-tuning. On LIBERO, we focus on comparing the performance with other methods after fine-tuning. All evaluation tasks follow [11, 12, 13, 27].
@@
Evaluation on SIMPLER  SIMPLER [32] is an open-source suite of simulated environments to evaluate policies. Visual Matching reduces the visual disparity by overlaying real-world backgrounds and adjusting object textures to align with real environments. Variant Aggregation introduces multiple randomized variations in the simulation’s visual elements. In SIMPLER, we test FPC-VLA on two commonly used robotic platforms: WidowX and Google Robot. Visual results can be found in Fig. 3.
WidowX is tested for grasping, stacking blocks, and placing items into containers. Each task is repeated 24 times under varying object poses. Table III compares various methods with FPC-VLA, reporting performance in success rates (%). RT-1-X [7] and OpenVLA [9] perform poorly across most tasks, with task success rates near 0%, indicating difficulty in generalization due to domain gaps. Notably, FPC-VLA even outperforms the fine-tuned SpatialVLA [10] under the zero-shot setting, demonstrating its effectiveness in both grasping and task completion across varied manipulation scenarios.
Google Robot evaluation includes pick-and-place operations and drawer opening/closing tasks. Each task is repeated under varying object poses and distractors. The number of attempts for each task is indicated in parentheses in Table IV. As shown in Table IV, in Visual Matching category, FPC-VLA achieves an average score of 78.0%, outperforming all baselines, despite in zero-shot setting. For Variant Aggregation tasks, which evaluates the ability to generalize across variants, FPC-VLA again ranks highest. It is worth noting that RT-1 [5] is trained solely on Google Fractal dataset [5], but FPC-VLA significantly outperforms it without any fine-tuning on that dataset. Even when faced with the most complex task (“Open Top Drawer and Place Apple” in Variant Aggregation setting), FPC-VLA still performs remarkably well, surpassing the results of other methods.
Evaluation on LIBERO   LIBERO [26] simulation platform features four distinct task suites: LIBERO-Goal concentrates on achieving different task goals with fixed objects and spatial arrangements. LIBERO-Spatial focuses on adapting to varying spatial configurations of objects while maintaining object consistency. LIBERO-Object involves interacting with a diverse set of objects. LIBERO-Long contains long-horizon tasks that involve diverse objects, layouts, and goals.
@@
TABLE VI: Results of ablation study
@@
Put spoon on towel
Put carrot on plate
Stack green block
on yellow block
@@
Avg. Grasp Avg. Success
Grasp Success Grasp Success Grasp Success Grasp Success
Supervisor w/o Supervisor 83.3 75.0 62.5 50.0 50.0 37.5 79.2 70.8 50.0 58.3
w/o correction 75.0 75.0 58.3 45.8 58.3 45.8 83.3 66.7 68.7 58.3
@@
29.2 12.5 41.7 8.3 41.7 20.8 16.7 0 32.3 10.4
directly average 79.2 79.2 58.3 54.2 62.5 50.0 83.3 66.7 70.8 62.5
w/o time decay 79.2 70.8 50.0 37.5 58.3 33.3 79.2 66.7 66.7 52.1
motion disturbance + w/o Supervisor 62.5 54.2 50.0 41.7 50.0 20.8 37.5 25.0 50.0 35.4
motion disturbance + FPC-VLA 75.0 62.5 62.5 54.2 62.5 37.5 62.5 50.0 65.6 51.1
FPC-VLA 79.2 79.2 58.3 58.3 70.8 45.8 79.2 75.0 71.9 64.6
Following the data modification and augmentation procedures from [9], we regenerate higher-resolution images (256×256px) , remove no-op actions, rotate third-person images by 180°, filter out failed demonstrations, and use only third-person camera views for fair comparison. Each task suite is evaluated over 500 trials. As shown in Table V, FPC-VLA achieves the highest average performance of 86.9%, significantly outperforming all other baselines. Notably, on the most challenging long-horizon tasks, it scores 82.2%, well above the second-best 70.9% by ThinkAct [27], indicating that failure prediction and correction can help avoid deviation in long-term tasks.
IV-C Real-World Experiments
To assess FPC-VLA’s real-world applicability, as shown in Fig. 4, we evaluate five challenging tasks on the Xiaomi Robot and ALOHA. Xiaomi Robot is a wheeled dual-arm system with 7-DoFs per arm, considering only the gripper’s open/close DoF. ALOHA has dual 6-DoF arms with parallel grippers. For each task, FPC-VLA is fine-tuned with 1000 in-domain demonstrations, while 1000 failure–correction pairs are used to fine-tune the supervisor. Inference uses FP16 on a single NVIDIA RTX 3090, with 32 trials per task. Results show that FPC-VLA achieves robust performance on long-horizon, cluttered tasks, with an average success rate of 86.3%. Notably, we include a dual-arm handover task (Put cup on the left of table), which demonstrates FPC-VLA’s versatility across multi-configuration robots.
IV-D Evaluation of Failure Prediction and Correction
To evaluate the accuracy of our supervisor’s failure prediction and correction, we compare it with several advanced VLMs [33, 34, 35, 31]. Our supervisor is fine-tuned using the failure correction dataset generated as described in Section III-C, while the other VLMs use publicly available pretrained weights. As shown in Fig. 5, our supervisor outperforms all other methods on failure correction datasets derived from three different robotics datasets. Notably, our approach surpasses the base model Qwen2.5-VL-7B using only fine-tuning on generated data.
IV-E Ablation Study

### v2

@@ 4 Experiment
RT-1 [brohan2022rt] 85.7 44.2 73.0 6.5 52.4 89.8 50.0 32.3 2.6 43.7
RT-1-X [o2024open] 56.7 31.7 59.7 21.3 42.4 49.0 32.3 29.4 10.1 30.2
RT-2-X [o2024open] 78.7 77.9 25.0 3.7 46.3 82.3 79.2 35.3 20.6 54.4
Octo-Base [team2024octo] 17.0 4.2 22.7 0.0 11.0 0.6 3.1 1.1 0.0 1.2
OpenVLA [kim2024openvla] 18.0 56.3 63.0 0.0 34.3 60.8 67.7 28.8 0.0 39.3
RoboVLMs [li2024towards] 77.3 61.7 43.5 24.1 41.9
CogACT [li2024cogact] 91.3 80.5 71.8 50.9 74.8 89.6 80.8 28.3 46.6 61.3
SpatialVLA [qu2025spatialvla]
@@
Table 5: Comparison results on LIBERO with Franka
@@
Diffusion [chi2023diffusion] 68.3 78.3 82.5 50.5 72.4
Octo-Base [team2024octo] 84.6 78.9 85.7 51.1 75.1
OpenVLA [kim2024openvla] 79.2 84.7 88.4 53.7 76.5
TravceVLA [zheng2024tracevla] 75.1 84.6 85.2 54.1 74.8
SpatialVLA [qu2025spatialvla] 78.6 88.2 89.9 55.5 78.1
ThinkAct [thinkact] 87.1 88.3 91.4 70.9 84.4
Cot-VLA [zhao2025cot] 87.6 87.5 91.6 69.0 81.1
GRAPE [zhang2024grape] 83.1 88.5 92.1 57.2 80.2
@@ 4.1 Implementation Details
The VLM model of the supervisor uses Qwen2.5-vl 7B [bai2025qwen2]. During training, we use 16 NVIDIA H100 GPUs. For the VLA model, the batch size is set to 256, the learning rate is $2\times 10^{-5}$, and 15 action steps are predicted per inference. VLA model is pretrained on the mixture of OXE dataset [o2024open] and LIBERO dataset [liu2023libero]. For the supervisor model, training is conducted with bfloat16 precision, a batch size of 128, a learning rate of $10^{-4}$, LoRA rank of 8, LoRA alpha of 32, and max pixels of 1003520. Synthetic failure dataset is generated from BridgeV2 [walke2023bridgedata] for WidowX, [brohan2022rt] for Google Robot, and [liu2023libero] for Franka. Each robot has a synthetic failure dataset containing 100k entries, while annotated failure dataset generated in MuJoCo and collected from the real robot via teleoperation contain 10k entries. In action fusion module, $\alpha=0.1$, $\lambda=\beta=0.01$. FPC-VLA achieves inference time of 0.176s for non-keyframes and 1.766s for keyframes. As the supervisor is called at most three times during the hundreds of inferences per task, its benefit becomes more evident in long-horizon tasks. For example, in the Open Top Drawer and Place Apple task, the total inference time of FPC-VLA is 40.5 s, which is only slightly higher than the 35.2 s when the supervisor is inactive.
4.2 Evaluation on Simulation Benchmarks
@@
To evaluate the FPC-VLA model, we conduct experiments on two simulation platforms—SIMPLER [xuan2024eval] and LIBERO [liu2023libero] —as well as on three types of robotic arms: WidowX, Google Robot, and Franka. We primarily highlight FPC-VLA’s zero-shot learning capability, without domain-specific fine-tuning. The terminology regarding zero-shot and fine-tuning here is consistent with that used in [qu2025spatialvla]. All evaluation tasks follow [zheng2024tracevla, li2024towards, li2024cogact, thinkact].
@@
Evaluation on SIMPLER   SIMPLER [xuan2024eval] is an open-source suite of simulated environments specifically designed for the comprehensive evaluation of robotic policies. To enhance visual realism and bridge the sim-to-real gap, SIMPLER incorporates two key techniques: Visual Matching and Variant Aggregation. Visual Matching reduces visual disparity between simulation and reality by overlaying real-world background images and fine-tuning object textures so they better align with those found in actual physical environments. Meanwhile, Variant Aggregation introduces multiple randomized variations across visual elements—such as lighting conditions, object appearances, and scene layouts—to increase the diversity and robustness of the training and evaluation settings. In SIMPLER, we test FPC-VLA on two commonly used robotic platforms: WidowX and Google Robot. Visual results can be found in Fig. 4.
@@
WidowX is evaluated on a suite of manipulation tasks, including grasping individual objects, stacking blocks, and placing items into containers. To assess robustness and generalization, each task is repeated 24 times under diverse initial object poses and environmental configurations. As summarized in Table 3, we compare the performance of several state-of-the-art methods against our proposed FPC-VLA framework, reporting results in terms of grasping success rates and task success rates (%). Notably, both RT-1-X [o2024open] and OpenVLA [kim2024openvla] exhibit poor performance across nearly all tasks, achieving success rates close to 0%. This suggests significant challenges in transferring policies trained in simulation or on different robot platforms to the WidowX setup, likely due to substantial domain gaps. In contrast, FPC-VLA demonstrates strong zero-shot generalization capabilities—remarkably outperforming even the fine-tuned SpatialVLA [qu2025spatialvla] without any task-specific adaptation, , demonstrating its effectiveness in both grasping and task completion across varied manipulation scenarios, demonstrating its effectiveness in both grasping and task completion across varied manipulation scenarios.
Google Robot evaluation includes pick-and-place operations and drawer opening/closing tasks. To rigorously assess generalization capabilities, each task is executed repeatedly under diverse conditions—specifically, with varying object poses and the presence of visual distractors. The total number of attempts per task is explicitly reported in parentheses in Table 4. As shown in Table 4, in Visual Matching category, FPC-VLA achieves an average score of 78.0%, outperforming all baselines, despite in zero-shot setting. For Variant Aggregation tasks, which evaluates the ability to generalize across variants, FPC-VLA again ranks highest among all compared methods. It is worth noting that RT-1 [brohan2022rt] is trained solely on Google Fractal dataset [brohan2022rt], but FPC-VLA significantly outperforms it without any fine-tuning on that dataset. Even when faced with the most complex task (“Open Top Drawer and Place Apple” in Variant Aggregation setting), FPC-VLA still performs remarkably well, surpassing the results of other methods.
Evaluation on LIBERO    LIBERO [liu2023libero] simulation platform features four distinct task suites: LIBERO-Goal focuses on achieving a variety of task objectives while keeping the set of objects and their spatial arrangements fixed, thereby isolating the challenge of goal-conditioned policy learning. LIBERO-Spatial emphasizes adaptation to varying spatial configurations of the same objects, testing the model’s ability to generalize across different layouts without changes in object identity. LIBERO-Object introduces diversity in the object set itself, requiring robots to interact with a wide range of previously unseen objects, thus evaluating object-level generalization. LIBERO-Long contains long-horizon tasks that involve diverse objects, layouts, and goals.
Following the data modification and augmentation procedures from [kim2024openvla], we regenerate higher-resolution images (256×256px) , remove no-op actions, rotate third-person images by 180°, filter out failed demonstrations, and use only third-person camera views for fair comparison. Each task suite is evaluated over 500 trials. As shown in Table 5, FPC-VLA achieves the highest average performance of 86.9%, significantly outperforming all other baselines despite not using in-domain data for fine-tuning. Notably, on the most challenging long-horizon tasks, it scores 82.2%, well above the second-best 70.9% by ThinkAct [thinkact], indicating that failure prediction and correction can help avoid deviation in long-term tasks.
4.3 Real-World Experiments
To assess FPC-VLA’s real-world applicability, as shown in Fig. 5, we evaluate five challenging tasks on the Xiaomi Robot and ALOHA.
Figure 6: Xiaomi Robot and ALOHA for real world experiments.
Table 6: Comparison results in real world experiments
Robot Xiaomi Robot ALOHA Robot Avg. Grasp Avg. Task
Method
Stack orange block
on green block
Put white ball
and green ball
in blue plate
Stack orange block
on yellow block
Put cup on
the left of table
Put yellow block
and red cylinder
in white plate
Grasp Task Grasp Task Grasp Task Grasp Task Grasp Task
ACT [zhao2023learningfinegrainedbimanualmanipulation] 62 34 58 52 64 30 72 58 52 46 61.6 44.0
Diffusion [chi2023diffusion] 44 12 22 18 14 8 36 26 38 32 30.8 19.2
OpenVLA [kim2024openvla] 42 22 38 36 54 28 56 42 28 24 43.6 30.4
$\pi_{0}$ [black2024pi0visionlanguageactionflowmodel] 58 42 60 54 72 32 44 36 40 34 54.8 39.6
$\pi_{0}$-Fast [pertsch2025fastefficientactiontokenization] 38 12 44 42 54 40 60 48 46 38 48.4 36.0
CogACT [li2024cogact] 82 70 86 82 88 78 90 82 84 80 86.0 78.4
FPC-VLA 96 84 88 88 94 82 96 90 90 86 92.8 86.0
Xiaomi Robot Setup   Fig.6(a) illustrates our 7-DoF Xiaomi Robot setup. The head is equipped with an Intel RealSense D435 RGB-D camera, which captures close-range task observations by streaming 640×480 RGB-D frames at 30 Hz. This configuration supports real-time workspace perception, enabling capabilities such as object detection and pose estimation. All policy outputs are provided as Cartesian coordinate increments expressed in the robot’s base frame, facilitating seamless integration with the arm’s control system through a ROS-based interface. Data collection occurs at 10 Hz via teleoperation, ensuring compatibility with low-latency applications.
ALOHA Robot Setup   Our ALOHA hardware platform (Fig. 6(b)) is built upon the Tsinghua Airbot system. It features two lightweight 6-degree-of-freedom (6-DoF) manipulators mounted on a compact mobile base. Each arm is fitted with a 1-DoF parallel gripper and a wrist-mounted RGB-D camera to support close-range manipulation perception. Additionally, an Intel RealSense D435 RGB-D camera is attached to an adjustable overhead mast, capturing RGB-D frames at a resolution of 640×480 and a rate of 30 Hz. Whenever the pose of the overhead camera changes, we perform hand-eye calibration using an AprilTag calibration board to estimate the rigid transformation between the robot base frame and the camera frame. During preprocessing, all teleoperated trajectories are transformed into the overhead camera’s coordinate frame. Conversely, actions generated by the learned policy are mapped back to the robot base frame using the calibrated extrinsic parameters. The model outputs Cartesian incremental commands, which are integrated over time to produce target end-effector poses. These poses are then converted into joint-space commands via a damped least-squares inverse kinematics solver, with explicit enforcement of joint velocity and acceleration limits.
Real World Data Preparation   For the Xiaomi Robot, we collect data by controlling the robotic arm using a SpaceMouse. For ALOHA, we use its built-in dual-arm teleoperation system to collect trajectories containing joint positions, end-effector poses, and RGB-D images. The raw data undergoes several preprocessing steps: first, we apply a median filter with a kernel size of 5 and remove outlier frames exceeding a threshold of three standard deviations to eliminate noise; second, we temporally synchronize the multimodal data streams and downsample the ALOHA data from 30 Hz to 10 Hz to match the temporal frequency of the Xiaomi robot data, ensuring temporal alignment across datasets; finally, we convert the processed data into the RLDS format, structuring it to align with the OXE dataset schema, which includes standardized fields for observations, actions, and rewards.
Analysis of Comparison Results   Table 6 presents a comprehensive comparison of various robotic manipulation methods across multiple real-world tasks performed on two distinct robotic platforms: the Xiaomi Robot and the ALOHA Robot. The evaluation metrics include grasp success rate (Grasp) and task completion success rate (Success). Inference uses FP16 on a single NVIDIA RTX 3090, with 50 trials per task. Our proposed method, FPC-VLA, consistently outperforms all baseline approaches across both robots and all individual tasks. While $\pi_{0}$ [45] and OpenVLA [15] show moderate performance, methods like Diffusion [40] and $\pi_{0}$-Fast [46] struggle with low success rates, especially on the ALOHA platform. ACT [44] performs reasonably well but is still substantially outperformed by FPC-VLA in both grasp precision and task completion reliability. FPC-VLA achieves the highest average grasp success rate (92.8%) and task success rate (86.0%), significantly surpassing the second-best method, CogACT [19], which attains 86.0% and 78.4%, respectively. As shown in Fig. 5, thanks to the VLM-based Supervisor, FPC-VLA demonstrates strong failure prediction and correction capabilities, effectively reducing instances of grasping or placement failures. Notably, we include a dual-arm handover task (Put cup on the left of table), which demonstrates FPC-VLA’s versatility across multi-configuration robots.
Figure 7: Evaluation of Supervisor cross-platform capability. Success rates (%) are reported for models trained on Google Fractal and BridgeV2 datasets, and evaluated on WidowX robot.
4.4 Evaluation of Cross-Platform Generalization Capability
To demonstrate the cross-platform generalization capability of our supervisor, we conduct experiments on the WidowX robot in SIMPLER using a supervisor fine-tuned exclusively on data from [brohan2022rt] (Google Robot). This setting evaluates whether the failure prediction and correction policies learned on one robotic platform can effectively transfer to another with different morphology, actuation, and visual appearance—without any additional fine-tuning or adaptation. The results, presented in Fig. 7, show that our supervisor achieves competitive success rates across all evaluated tasks on the WidowX robot, despite being trained solely on Google Robot data. The remarkable zero-shot performance arises from several synergistic design principles in FPC-VLA: First, the supervisor reasons over high-level, task-grounded natural language queries instead of low-level robot states, decoupling failure analysis from embodiment details. Second, corrections use a standardized, egocentric vocabulary of relative directions and magnitudes, enabling platform-agnostic transferability compared to absolute action deltas. Thirdly, the failure-correction dataset encodes only task-relevant spatial changes into language, capturing universal manipulation heuristics that generalize across robot embodiments.
4.5 Evaluation of Failure Prediction and Correction
To evaluate the accuracy of our supervisor’s failure prediction and correction capability, we conduct a comparative study against several state-of-the-art vision–language models (VLMs), including DeepSeek-VL-7B [zhu2024deepseek], InternVL3-8B [zhu2025internvl3], LLaVA-1.5-7B [liu2024improved], and the base Qwen2.5-VL-7B [bai2025qwen2]. Following standard practice for evaluating task-specific adaptation, our supervisor is fine-tuned on the automatically generated failure correction dataset described in Section 3.4, which provides structured vision–language–action supervision tailored to robotic manipulation failures. In contrast, the baseline VLMs are evaluated using their publicly released, general-purpose pretrained weights without any task-specific fine-tuning or prompt engineering beyond what is necessary to align with our input format.
Table 7: Results of ablation study
Robot WidowX Xiaomi robot ALOHA Avg. Grasp Avg. Task
@@ 4.5 Evaluation of Failure Prediction and Correction
Put spoon
on towel
@@
Stack orange block
on green block
Put cup on
the left of table
Grasp Task Grasp Task Grasp Task Grasp Task
Supervisor w/o Supervisor 70.8 70.8 79.2 70.8 84 72 90 84 81.0 74.4
w/o correction 75.0 70.8 83.3 70.8 84 74 92 86 83.6 75.4
@@
29.2 12.5 41.7 33.3 52 40 72 66 48.7 38.0
directly average 79.2 70.8 83.3 66.7 84 76 86 82 83.1 73.9
w/o time decay 79.2 79.2 79.2 66.7 92 82 94 90 86.1 79.5
motion disturbance + w/o Supervisor 62.5 54.2 37.5 25.0 72 60 72 64 61.0 50.8
motion disturbance + FPC-VLA 75.0 62.5 62.5 50.0 80 74 82 78 74.9 66.1
FPC-VLA 79.2 79.2 79.2 75.0 96 84 96 90 87.6 82.1
As shown in Fig. 8, our fine-tuned supervisor consistently outperforms all baselines on WidowX, Google Robot and Franka. This performance gap highlights the effectiveness of both our data generation pipeline and the alignment between the supervision signal and the target task. Importantly, even when compared to its own base architecture Qwen2.5-VL-7B, our approach achieves a significant improvement solely through fine-tuning on synthetically generated, robotically grounded QA pairs. Our results suggest that while general-purpose VLMs possess strong visual and linguistic capabilities, they benefit substantially from targeted fine-tuning on semantically structured, robot-relevant data—even if synthesized—to achieve reliable performance in safety-critical decision-making contexts.
4.6 Ablation Study